OpenStudy (anonymous):

the main frame entrance to the louvre museum in oaris, france, is a square pyramid made entirely of glass. one side of the base is 35 meters and the hieght is 20,6 meter. What is the surface area of the louvre pyramid. plz show work

OpenStudy (anonymous):

Area of Triangle = (1/2)*Base*Height Area of one side of the Lourve = (1/2)*35*20.6 = 360.5 m^2 Area of 4 sides of the Lourve = 360.5 * 4 = 1442 m^2 I am assuming the square base of the Lourve is not being considered for the surface area.